# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

### v1

@@ 5 Proofs of Main Results
Similarly to Lemma 4.5, the fact that $k_{t}$ increases with the number of outer-loop iterations $t$ implies that the errors introduced by the Richardson iteration method decrease with increasing $t$. However, in contrast with Lemma 4.5 where $\mathbf{D}^{(t-1)}$ is fixed, the dependence of $\mathbf{D}^{(t-1)}$ on $\bm{\pi}^{(t-1)}$ implies that the right-hand side of the linear system solutions in Step $\mathbf{5}$ of Algorithm 2 is also perturbed. Nonetheless, the operations involved in $\mathbf{D}^{(t-1)}$ are linear and thus the perturbation introduced at iteration $t$ are of the same order as the error introduced by the Richardson iteration method in the previous iteration $t-1$. Hence, as long as the application of the Richardson iteration method becomes more accurate over time, Algorithm 2 converges to the true stationary distribution $\bm{\pi}$ analogous to the arguments establishing Theorem 4.2.
More specifically, the update at the outer-loop iteration $t$ satisfies
$\overline{\bm{\pi}}^{(t)\top}=(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top}+(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\overline{\mathbf{D}}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top},$
where $\overline{\mathbf{D}}^{(t-1)}$ denotes the part of $\mathbf{D}^{(t-1)}$ attributed to the non-exact solution of the linear system in the previous iteration $t-1$ and is of the order $O\left(\|\mathbf{D}^{-\top}\|\dfrac{\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|^{k_{t-1}+1}}{1-\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|}\right)$. Ignoring the second term which depends quadratically on the error tolerance used in the Richardson iteration method, we are left with
$\overline{\bm{\pi}}^{(t)\top}$ $\approx(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top}$
$=\left(\mathbf{D}^{\top}-\mathbf{U}^{\top}\right)^{-1}-\left((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots\right)\mathbf{D}^{-\top}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top},$
which eventually reduces to the iterative behavior of Algorithm 1 and thus the convergence follows from the arguments establishing Theorem 4.2. This yields the desired result. ∎
@@ 6 Numerical Experiments
In this section we present a representative sample of numerical experiments that support our theoretical results and empirically evaluate our general classes of algorithmic approaches in comparison with the most efficient existing numerical methods for solving NCD Markov processes. Since the most efficient existing algorithms are the iterative numerical methods due to Takahashi [27], Vantilborgh [28], and Koury, McAllister and Stewart [18], with the latter being the most recent of the three methods, and since all three numerical methods exploit aggregation-disaggregation in a very similar manner with similar convergence behaviors [7], we selected the Koury-McAllister-Stewart (KMS) algorithm as the baseline method for the numerical evaluation comparison of our algorithms.
We first present details on the set up of our numerical experiments and then we present and discuss the performance results from a representative sample of our numerical experiments. Beyond quantitatively supporting our theoretical results, these empirical performance results demonstrate that our general classes of algorithmic approaches exhibit relatively little or no increase in the number of outer-loop iterations and orders of magnitude improvements in the computation time, both in comparison with the baseline method of the original KMS algorithm.
@@ 6.1 Experimental Setup
Our numerical experiments of solving NCD Markov processes with Algorithm 1, Algorithm 2 and the baseline original KMS algorithm are conducted in Python on a standard compute infrastructure with a $2.5$Ghz i7777 processor and $100$GB of RAM. The $n\times n$ transition probability matrix $\mathbf{P}$ of block structural form (2) for each independent trial run was randomly generated with each element independently and identically sampled from a uniform distribution on the unit interval. Then, to satisfy the conditions (3), we reduce the magnitude of the elements of the off-diagonal block matrices by proportionally scaling these elements such $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m]$, $i\neq j$; and, to satisfy the requirement that $\mathbf{P}$ be a stochastic matrix, we correspondingly and proportionally scale each row of the matrices $\mathbf{P}_{ii}$, $i\in[m]$, so that each row of the resulting matrix $\mathbf{P}$ sum to one.
@@
Next, we execute every step of each of the three algorithms on the above computational infrastructure, computing the number of outer-loop iterations and computing the number of operations for each step, with the latter converted to obtain the computation time. For every algorithm, the number of operations required to compute the solution of a $k\times k$ linear system using LU decomposition is given by $3k^{3}$, with the operations per second depending upon the precision used for the computation. Other than Step $\mathbf{5}$ and Step $\mathbf{3}$ which are performed in lower precision in Algorithm 1 and Algorithm 2 (see below), all three algorithms execute the same number of operations for each outer-loop iteration which are performed in full-precision. The computation time for each aspect of each algorithm is then obtained by multiplying the number of operations by the processor performance for the level of precision used for the computation. For this purpose, we used the performance of the NVIDIA H100 processor [23, 2] whose peak performance for double-, single- and half-precision floating point arithmetic are 34, 67 and 134 Tflops per second, respectively. The baseline original KMS algorithm performs all its computations in double-precision; Algorithm 1 performs the computations of Step $\mathbf{5}$ and Step $\mathbf{3}$ typically in single-precision; Algorithm 2 performs the computations of Step $\mathbf{5}$ in single-precision and the computations of Step $\mathbf{3}$ in full-precision.
@@
The above computation times are only used when the matrix of each linear system fits within the available memory on the processor architecture. In particular, when the size of the matrix $\mathbf{P}_{ii}$ at the level of precision employed exceeds the available memory of $96$GB, then the computation of the linear system becomes memory bandwidth limited as described in Section 4.3. Based on the specification for the NVIDIA H100 processor [23, 2], the memory bandwidth of $3.35$ Tbytes per second is instead used when any of the algorithms exceed the available memory and become bandwidth limited. As noted in Section 4.3, the use of lower precision computation in our general algorithmic approaches makes it possible to handle larger systems of linear equations before hitting the memory bandwidth limitations of the processor architecture.
@@
Lastly, it is important to note that the performance benefits of Algorithm 1 and Algorithm 2 over the baseline original KMS algorithm stem from a combination of the higher performance throughput of the lower precision arithmetic and the reduced operations of the iterative approximate computing methods.
@@ 6.2 Empirical Performance Results
We now present the performance results from our numerical experiments of solving NCD Markov processes with Algorithm 1 and Algorithm 2 that support our theoretical results and quantify the empirical benefits of our general algorithmic approaches over the baseline original KMS algorithm. The primary parameters that influence the computation of the solution of (1) to obtain the stationary distribution $\bm{\pi}$ of NCD Markov processes concern: ($a$) the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$; ($b$) the number $m$ of diagonal block matrices; and ($c$) the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. We consider collections of performance result comparisons from numerical experiments in which two of the primary parameters are fixed and the remaining primary parameter varies over a range of values. Each set of performance result comparisons comprises evaluation of our conservatively modified KMS approach — Algorithm 1 — denoted by KMS+IR, our aggressively modified KMS approach — Algorithm 2 — denoted by KMS+Ri, and the baseline KMS algorithm. The quantitative metrics of interest in each collection of performance comparisons are the mean number of iterations and the mean computation time in seconds, taken over ten independent experimental trial runs.
@@
For Algorithm 1, we select the computational precision level with respect to the condition number and the norm of the matrix of the linear systems of Step $\mathbf{5}$ and Step $\mathbf{3}$ as defined in Section 3.1. We then employ the iterative refinement method as described in Section 3.1 to compute the solution of each linear system to full-precision accuracy. For Algorithm 2, we employ the Richardson iteration method with block-Jacobi preconditioning as described in Section 3.2 to solve the linear systems of Step $\mathbf{5}$ with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as the iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as the right-hand side. In particular, the linear systems of Step $\mathbf{5}$ with the coefficient matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ are preconditioned by the block-diagonal matrix $\mathbf{D}^{\top}$, where the preconditioner is applied via LU factorization computed in $32$-bit precision. For the solution of each linear system of Step $\mathbf{5}$, we start with a number of maximum-allowed iterations equal to ten and increase this value by five per outer iteration. The linear system of Step $\mathbf{3}$ is solved exactly in full precision so that we can focus on the performance benefits of our algorithmic approach for the larger linear systems of Step $\mathbf{5}$.
We present the results from a representative sample of our numerical experiments. Figure 1 presents the set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $200$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the number $m$ of diagonal block matrices is varied such that $m\in\{5,10,20,50,100\}$. We first observe from Figure 1(a) that the mean number of iterations under KMS+IR are either identical to that of the original KMS algorithm or differ by at most one, with the mean number of iterations under KMS+Ri only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains fairly flat for all algorithms, which is as expected since the linear increase is in the number of linear systems of the same size being solved. From Figure 1(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time and that such performance improvements increase with the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation times for each algorithm increase with $m$ as expected due to the linear increase in the number of linear systems of the same size being solved, with the rate of increase greatest for the original KMS algorithm, with much smaller rates of increase for both of our general algorithmic approaches, and with even smaller rates of increase for KMS+Ri than for KMS+IR. We note that our choice for the number of maximum-allowed iterations in the solution of each linear system of Step $\mathbf{5}$ under KMS+Ri leads to the lowest mean computation times in Figure 1(b). The increase in the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, under KMS+Ri increases the part of the iteration matrix not captured by the block-Jacobi preconditioner, but this use of the method becomes computationally less expensive and thus there is an interesting algorithmic tradeoff which can be exploited within the context of specific applications. Meanwhile, the rate of increase under KMS+Ri as a function of the number of block matrices is the smallest among all algorithms.
@@
Figure 1: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for diagonal matrices $\mathbf{P}_{ii}$ with dimension $n_{i}=200$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ with $\epsilon=0.1$, as a function of the number of blocks $m$. Performance results represent averages over $10$ independent trial runs.
Figure 2 presents the set of performance result comparisons for the case where the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are varied such that $n_{i}\in\{100,200,500,1000\}$. We first observe from Figure 2(a) that, once again, the mean number of iterations under KMS+IR are essentially identical to that of the original KMS algorithm, with the mean number of iterations under KMS+Ri only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains relatively flat for all algorithms, which is as expected due to the shared properties of the outer loop of all three algorithms. From Figure 2(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time and that such performance improvements are consistent and increasing across the various dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation times for each algorithm increase with $n_{i}$ as expected due to the linear increase in the size of the linear systems being solved, with the rate of increase somewhat larger for the original KMS algorithm, with somewhat smaller rates of increase for both of our general algorithmic approaches, and with somewhat smaller rates of increase for KMS+Ri than for KMS+IR. Once again, we note that our general KMS+Ri approach provides the lowest mean computation times for the reasons described above, although the differences with our general KMS+IR approach is relatively small but grows somewhat with increases in the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$.
@@
Figure 2: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for number of blocks $m=20$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ with $\epsilon=0.1$, as a function of the dimension $n_{i}$ of the diagonal matrices $\mathbf{P}_{ii}$. Performance results represent averages over $10$ independent runs.
Figure 3 presents the set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $200$, the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, and the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is varied such that $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$. We first observe from Figure 3(a) that the mean number of iterations under KMS+IR are identical to that of the original KMS algorithm, with the mean number of iterations under KMS+Ri somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, remains relatively flat for both KMS and KMS+IR, as expected again due to the shared properties of the outer loop of all three algorithms. However, as observed in Figure 3(a), the mean number of iterations under KMS+Ri increases and requires more outer-loop iterations as $\epsilon$ increases. Due to the nature of the block-Jacobi preconditioner, the convergence of the Richardson iteration method becomes slower with increasing $\epsilon$ because now the part of the iteration matrix that is left outside the block-diagonal structure becomes more computationally significant. From Figure 3(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time with such performance improvements remaining flat across the various magnitudes $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. The mean computation times for each algorithm remain relatively flat as a function of the magnitude $\epsilon$, which is as expected due to the shared properties of the outer loop of all three algorithms. We note that our general KMS+Ri approach provides somewhat lower mean computations times than KMS+IR for smaller values of $\epsilon$ and that KMS+IR provides somewhat lower mean computations times than KMS+Ri for larger values of $\epsilon$. This is due to the nature of the block-Jacobi preconditioner, as described above.

### v2

@@ 5.1 Proof of Lemma 17
Let $F$ denote the (fixed) process defined by a single outer iteration of Algorithm 2, and thus ${\bm{\pi}}^{(t)\top}=F({\bm{\pi}}^{(t-1)\top})$. By supposition, we have
$\|{\bm{\pi}}^{(t-1)\top}-{\bm{\pi}}^{\top}\|=O(\varphi).$ (33)
Now, instead of ${\bm{\pi}}^{(t)\top}$, assume that Algorithm 2 computes
${\bm{\pi}}^{(t)\top}=F({\bm{\pi}}^{(t-1)\top})+\mathbf{r}_{t}^{k_{t}},$ (34)
where $\mathbf{r}_{t}^{k_{t}}$ denotes the approximation error resulting from the application of $k_{t}$ steps of the RI method. Subtracting ${\bm{\pi}}^{\top}$ from both sides of (34), we have
${\bm{\pi}}^{(t)\top}-{\bm{\pi}}^{\top}=F({\bm{\pi}}^{(t-1)\top})-{\bm{\pi}}^{\top}+\mathbf{r}_{t}^{k_{t}},$
and thus taking the norms of both sides and substituting for $\mathbf{r}_{t}^{k_{t}}$ yields
$\|{\bm{\pi}}^{(t)\top}-{\bm{\pi}}^{\top}\|\leq\|F({\bm{\pi}}^{(t-1)\top})-{\bm{\pi}}^{\top}\|+\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k_{t}}\|\mathbf{L}^{\top}\mathbf{D}^{(t-1)}{\bm{\pi}}^{(t-1)\top}\|.$
It then follows from analogous arguments establishing Theorem 4.2 that $\|F({\bm{\pi}}^{(t-1)\top})-{\bm{\pi}}^{\top}\|=O(\epsilon)\|{\bm{\pi}}^{(t-1)\top}-{\bm{\pi}}^{\top}\|$, and therefore we can simplify the above equation as
$\|{\bm{\pi}}^{(t)\top}-{\bm{\pi}}^{\top}\|\leq O(\epsilon)\|{\bm{\pi}}^{(t-1)\top}-{\bm{\pi}}^{\top}\|+\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k_{t}}\|\mathbf{L}^{\top}\mathbf{D}^{(t-1)}{\bm{\pi}}^{(t-1)\top}\|.$
Recall though that $k_{t}$ is always set so that $\|\mathbf{r}_{t}^{k_{t}}\|\leq O(\varphi\epsilon)$. Therefore, we can replace $\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k_{t}}\|\mathbf{L}^{\top}\mathbf{D}^{(t-1)}{\bm{\pi}}^{(t-1)\top}\|$ by its upper-bound, leading to
$\|{\bm{\pi}}^{(t)\top}-{\bm{\pi}}^{\top}\|$ $\leq O(\epsilon)\|{\bm{\pi}}^{(t-1)\top}-{\bm{\pi}}^{\top}\|+O(\varphi\epsilon)$
$\leq O(\varphi\epsilon),$
where the last result follows from (33) and $\epsilon<1$. Finally, the fact that RI converges linearly with respect to $\rho(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})$ follows by recalling that $\mathbf{r}_{t}^{k_{t}}=(\mathbf{I}-\mathbf{D}^{\top}+\mathbf{U}^{\top})^{k_{t}}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}{\bm{\pi}}^{(t-1)\top}$. ∎
@@ 6 Numerical Experiments
In this section we present a representative sample of numerical experiments that support our theoretical results and empirically evaluate our general mathematical framework of numerical methods in comparison with the most-efficient existing algorithms for computing the stationary distribution of NCD Markov processes. The most-efficient existing algorithms are the iterative numerical methods due to Takahashi [13], Vantilborgh [14], and Koury, McAllister and Stewart [15] (KMS), all three of which exploit aggregation-disaggregation in a very similar manner with similar convergence behaviors [17]. Although our general algorithmic framework of computational approaches can be applied within the context of any of these existing numerical methods and beyond, we again select the KMS algorithm as the appropriate baseline method for the numerical evaluation comparison of the two representative examples of our algorithmic framework in Section 3, since KMS is the most recent of the most-efficient existing algorithms.
We first present details on the set up of our numerical experiments and then we present and discuss the performance results from a representative sample of our numerical experiments. Beyond quantitatively supporting our theoretical results, these empirical performance results demonstrate that the two representative examples of our general algorithmic framework exhibit relatively little or no increase in the number of outer-loop iterations and orders of magnitude improvements in the computation time, both in comparison with the baseline method of the KMS algorithm.
@@ 6.1 Experimental Setup
Our numerical experiments of computing the stationary distribution of NCD Markov processes with Algorithm 1, Algorithm 2 and the baseline KMS algorithm are conducted in Python on a standard compute infrastructure with a $2.5$Ghz i7777 processor and $100$GB of RAM. For one primary set of numerical experiments, the $n\times n$ transition probability matrix $\mathbf{P}$ of block structural form (2) for each independent trial run is randomly generated with each element sampled independently and identically from a uniform distribution on the unit interval. Then, to satisfy the conditions (3), we reduce the magnitude of the elements of the off-diagonal block matrices by proportionally scaling these elements such that $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m]$, $i\neq j$; and, to satisfy the requirement that $\mathbf{P}$ be a stochastic matrix, we correspondingly and proportionally scale each row of the matrices $\mathbf{P}_{ii}$, $i\in[m]$, so that each row of the resulting matrix $\mathbf{P}$ sum to one. This first collection of numerical experiments essentially supports our empirical performance evaluation of Algorithm 1 and Algorithm 2 with respect to the various parameters of the NCD Markov processes, such as $m$, $\epsilon$ and $n_{i}$, $i\in[m]$. For another primary set of numerical experiments, we take a similar approach while exploiting a standard library of matrices for real-world linear systems [33]. In particular, as described above for each independent trial run, the $n\times n$ transition probability matrix $\mathbf{P}$ is randomly generated with each element sampled independently and identically from a uniform distribution on the unit interval. Then, the $n_{i}\times n_{i}$ diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are replaced by real-world matrices selected from the standard library in [33]. Next, to satisfy the conditions (3), we reduce the magnitude of the elements of the off-diagonal block matrices by proportionally scaling these elements such that $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m]$, $i\neq j$; and, to satisfy the requirement that $\mathbf{P}$ be a stochastic matrix, we correspondingly and proportionally scale each row of the matrices $\mathbf{P}_{ii}$, $i\in[m]$, so that each row of the resulting matrix $\mathbf{P}$ sum to one. This second collection of numerical experiments based on real-world matrices essentially supports and validates our first collection of numerical experiments based on randomly generated matrices, both with respect to the empirical performance evaluation of our Algorithm 1 and Algorithm 2.
@@
The performance metrics of interest taken over the independent trial runs of our numerical experiments are the mean number of outer-loop iterations and the mean computation time. To support some of the most advanced GPUs currently available in the marketplace, we developed a simulation of the computation times on a given GPU based on the specifications of the corresponding GPU architecture; and we validated our simulation of the computation times against physical experiments on the NVIDIA A100 processor [34]. More specifically, the computation times for every step of each of the three algorithms are simulated with respect to the number of operations for each step of the corresponding iterative process at its level of precision and the processor performance specification of the GPU of interest. For every algorithm, the computation times required to solve a $k\times k$ linear system are based on the use of LU decomposition at the precision level employed by the given algorithm for this linear system computation. Other than Step $\mathbf{5}$ and Step $\mathbf{3}$, which are performed in lower precision within Algorithm 1 and Algorithm 2 (see below), the simulated computation times for each outer-loop iteration are exactly the same for all three algorithms and are performed in full-precision. For the purpose of our numerical experiments, we used the performance of the NVIDIA H100 processor [30, 31] whose peak performance for double-, single- and half-precision floating point arithmetic are 34, 67 and 134 Tflops per second, respectively. The baseline KMS algorithm performs all of its computations in double-precision; Algorithm 1 performs the computations of Step $\mathbf{5}$ and Step $\mathbf{3}$ based on its precision-selection process, though typically in single-precision; Algorithm 2 performs the computations of Step $\mathbf{5}$ in single-precision and the computations of Step $\mathbf{3}$ in full-precision.
@@
The simulation of computation times described above are only used when the matrix of each linear system fits within the available memory on the GPU processor architecture. In particular, when the size of the matrix $\mathbf{P}_{ii}$, $i\in[m]$, at the level of precision employed exceeds the available memory of $96$GB, then the computation of the linear system becomes memory bandwidth limited as described in Section 4.3. Based on the specification for the NVIDIA H100 processor [30, 31], the memory bandwidth of $3.35$ Tbytes per second is instead used when any of the algorithms exceed the available memory and become bandwidth limited. As noted in Section 4.3, the use of lower-precision computation in our general algorithmic approaches makes it possible to handle larger systems of linear equations before hitting the memory-bandwidth limitations of the processor architecture.
@@
Lastly, it is important to note that the performance benefits of Algorithm 1 and Algorithm 2 over the baseline KMS algorithm in sharply reducing the core computational bottlenecks associated with solving systems of linear equations stem from a combination of the higher performance throughput of the lower-precision arithmetic and the reduced computations of the iterative approximate numerical methods.
@@ 6.2 Empirical Performance Results
We now present the performance results from our numerical experiments of computing the stationary distribution of NCD Markov processes with Algorithm 1 and Algorithm 2 that support our theoretical results and quantify the empirical benefits of our general mathematical framework of numerical methods over the baseline KMS algorithm. The primary parameters that influence the computation of the solution of (1) to obtain the stationary distribution $\bm{\pi}$ of NCD Markov processes concern: ($a$) the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$; ($b$) the number $m$ of diagonal block matrices; and ($c$) the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. We first consider collections of performance result comparisons from numerical experiments in which two of the primary parameters are fixed and the remaining primary parameter varies over a range of values. Each set of performance result comparisons comprises evaluation of the representative conservative example of our general algorithmic framework in Algorithm 1, the representative aggressive example of our general algorithmic framework in Algorithm 2, and the baseline KMS algorithm. As previously noted, the quantitative metrics of interest in each collection of performance comparisons are the mean number of iterations and the mean computation time in seconds, taken over ten independent experimental trial runs.
@@
For Algorithm 1, we select the computational precision level with respect to the condition number and the norm of the matrix of the linear systems of Step $\mathbf{5}$ and Step $\mathbf{3}$ as defined in Section 3.2. We then employ the IR method as described in Section 3.2 to compute the solution of each linear system to full-precision accuracy. For Algorithm 2, we employ the RI method with block-Jacobi preconditioning as described in Section 3.3 to solve the linear systems of Step $\mathbf{5}$ with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as the iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as the right-hand side. In particular, the linear systems of Step $\mathbf{5}$ with the coefficient matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ are preconditioned by the block-diagonal matrix $\mathbf{D}^{\top}$, where the preconditioner is applied via LU factorization computed in $32$-bit precision. For the solution of each linear system of Step $\mathbf{5}$, we start with a number of maximum-allowed iterations equal to ten and then geometrically increase this value per each outer iteration. The linear system of Step $\mathbf{3}$ is solved exactly in full precision so that we can focus on the performance benefits of our general algorithmic framework for the larger linear systems of Step $\mathbf{5}$.
@@
Figure 1: Performance comparison of our Algorithm 1 and Algorithm 2 with the KMS baseline for diagonal matrices $\mathbf{P}_{ii}$ having dimension $n_{i}=500$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ having $\epsilon=0.1$, as a function of the number of blocks $m\in\{5,10,20,50,100\}$. The comparative performance results represent averages over $10$ independent trial runs executed on a simulated H100100100100 GPU architecture.
We present in Figures 1 – 3 the results from a representative sample of our first collection of numerical experiments based on randomly generated $n\times n$ transition probability matrices $\mathbf{P}$ under different NCD process parameters. Figure 1 presents the representative set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $500$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the number $m$ of diagonal block matrices is varied such that $m\in\{5,10,20,50,100\}$. We first observe from Figure 1(a) that the mean number of iterations under Algorithm 1 are either identical to that of the baseline KMS algorithm or differ by at most one, with the mean number of iterations under Algorithm 2 only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains fairly flat for all algorithms, which is as expected since the linear increase is in the number of linear systems of the same size being solved. From Figure 1(b), we observe that both of the representative examples of our general algorithmic framework provide orders of magnitude reduction in the mean computation time and that such performance improvements increase with the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation time for each algorithm increases with $m$ as expected due to the linear increase in the number of linear systems of the same size being solved, with the rate of increase greatest for the baseline KMS algorithm, with much smaller rates of increase for both of the representative examples of our algorithmic framework, and with even smaller rates of increase for Algorithm 2 than for Algorithm 1. We note that our choice for the number of maximum-allowed iterations in the solution of each linear system of Step $\mathbf{5}$ under Algorithm 2 leads to the lowest mean computation times in Figure 1(b). The increase in the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, under Algorithm 2 increases the part of the iteration matrix not captured by the block-Jacobi preconditioner, but this use of the method becomes computationally less expensive and thus there is an interesting algorithmic tradeoff in Algorithm 2 which can be exploited within the context of specific applications. Meanwhile, the rate of increase under Algorithm 2 as a function of the number of block matrices is the smallest among all algorithms.
@@
Figure 2: Performance comparison of our Algorithm 1 and Algorithm 2 with the KMS baseline for number of blocks $m=20$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ having $\epsilon=0.1$, as a function of the dimension $n_{i}\in\{100,200,500,1000,2000\}$ of the diagonal matrices $\mathbf{P}_{ii}$. The comparative performance results represent averages over $10$ independent trial runs executed on a simulated H100100100100 GPU architecture.
Figure 2 presents the representative set of performance result comparisons for the case where the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are varied such that $n_{i}\in\{100,200,500,1000,2000\}$. We first observe from Figure 2(a) that, once again, the mean number of iterations under Algorithm 1 are essentially identical to that of the baseline KMS algorithm, with the mean number of iterations under Algorithm 2 only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains relatively flat for all algorithms, which is as expected due to the shared properties of the outer loop of all three algorithms. From Figure 2(b), we observe that both of the representative examples of our general algorithmic framework provide orders of magnitude reduction in the mean computation time and that such performance improvements are consistent and increasing across the various dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation time for each algorithm increases with $n_{i}$ as expected due to the linear increase in the size of the linear systems being solved, with the rate of increase somewhat larger for the baseline KMS algorithm, with somewhat smaller rates of increase for both of the representative examples of our algorithmic framework, and with somewhat smaller rates of increase for Algorithm 2 than for Algorithm 1. Once again, we note that our general approach in Algorithm 2 provides the lowest mean computation times for the reasons described above, although the differences in mean computation times with our general approach in Algorithm 1 are relatively small but tend to grow somewhat with increases in the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, with some differences from this trend due to the nature of the block-Jacobi preconditioner as described below.